If David has 19 bills in his wallet worth $125, all fives and tens, how many of each bill does he have?
Answer by checkley71(8403)

F+T=19 F=19-T
5F+10T=125 NOW SUBSTITUTE (19-T) FOR F IN THIS EQUATION & SOLVE FOR T
5(19-T)+10T=125
95-5T+10T=125
5T=125-95
5T=30
T=30/5
T=6 NUMBER OF TEN DOLLAR BILLS.
F=19-6
F=13 NUMBER OF FIVE DOLLAR BILLS.
PROOF
5*13+10*6=125
65+60=125
125=125